OpenStudy (anonymous):

Given MATH ~ ROKS, calculate the value of OK. A. 18 B. 17 C. 21 D. 20

OpenStudy (anonymous):

Information missing.

OpenStudy (anonymous):

Give me a moment.

OpenStudy (anonymous):

What does this question fall under? Category wise.

OpenStudy (lukecrayonz):

Since you know the two slanted sides are congruent, you know that the ratio is 5:25=1:5, so when given 4, multiply it by the ratio, 1:5, the answer is 20.

OpenStudy (anonymous):

I was trying to figure it out a different way.

Directrix (directrix):

Segments MA and RO are corresponding sides of similar quadrilaterals as are AT and OK. By definition of similar figures, corresponding sides of similar polygons are in proportion. MA is to RO as AT is to OK MA/RO = AT/OK 5/25 = 4/OK OK = ? @dopegyal_

OpenStudy (anonymous):

20

Directrix (directrix):

That is correct.

Directrix (directrix):

@xjakester >>What does this question fall under? Category wise. Geometry -- Plane Geometry -- Similar Polygons

OpenStudy (anonymous):

I knew it was geometry, I just couldn't remember the shape. I know I learned a different way to solve this, and I'm still going at it. @Directrix